2 The annual return on a stock is normally distributed with a mean of 10% and standard ation of 20%, with 95% confidence, we should expect its actual return in any par- devi ticular year to be between which pair of values? (Hint: Look again at page 12 of Lecture Note 1.) a.-5% and 35% b.-30% and 50% c.-45% and 75% d.-50% and 80% The margin requirement on a stock purchase is 20%. You fully use the margin allowed to loss? purchase 100 shares of AAPL at $150. If the price drops to $140, what is your percentage a. -20% b.-25% c. -33.3% d.-40% 3
Answer #1

Answer 2:

Correct answer is:

b. -30% and 50%

Explanation:

For data set following normal distribution, at 95% confidence level, values within two standard deviations of the mean account for approximately 95%.

At 95% confidence level the expected actual return will be between 'mean - 2 * Standard deviation' and 'mean + 2 * Standard deviation'

Expected actual return will be between (10% - 2 * 20% = ) -30% and (10% + 2 * 20% =) 50%.

As such option b is correct and other options a, c and d are incorrect.

Answer 3:

Correct answer is:

c. -33.3%

Explanation:

Purchase value of shares = 100 * $150 = $15,000

Amount invested = Margin requirement =$15,000 * 20% = $3,000

Price drops to $140:

Amount of loss = ($140 - $150) * 100 = - $1,000

% Loss = Loss amount / Amount invested = -1000 / 3000 = - 33.3%

As such option c is correct and other options a, b and d are incorrect.